OpenStudy (anonymous):

Please Help!! The number of items on two grocery lists differs by 7. The total number of items is 33. How many items are on each list?

OpenStudy (anonymous):

Hello! How far have you got with this question, so far?

OpenStudy (anonymous):

umm no where? i just started it

OpenStudy (anonymous):

i dont know where to start or end

OpenStudy (anonymous):

Okay, that's fine. Let's start by choosing a couple of variables (maybe call them x and y, or A and B). Any idea what they might represent?

OpenStudy (anonymous):

(It's okay if you have no idea!)

OpenStudy (anonymous):

okay lets do a and b.... and they represent the grocery list right?

OpenStudy (anonymous):

That's right, "a" would be the number of items on one list, and "b" would be the number of items on the other list.

OpenStudy (anonymous):

alright then what?

OpenStudy (anonymous):

Let's try writing down an equation (with a and b) to show that the total number of items is 33. Any idea how we might do that?

OpenStudy (anonymous):

(It's okay if you don't know!)

OpenStudy (anonymous):

well thats kinda my problem i have trouble trying to come up with the equatiion... i dont know

OpenStudy (anonymous):

Okay, no problem! The number of items in the first list is "a", and the number of items in the second list is "b". If the total number of items is 33, that means that adding up "a" and "b" will give you 33. Does that help?

OpenStudy (anonymous):

so your saying a plus b equals 33??

OpenStudy (anonymous):

Exactly! That's our first equation: a + b = 33.

OpenStudy (anonymous):

Let's be clear: "a" isn't the name of the first list. It's a number: the number of items in the first list (e.g. 10).

OpenStudy (anonymous):

Does this make sense, so far?

OpenStudy (anonymous):

ohhh okay i think so

OpenStudy (anonymous):

Great! Let's try the second clue then: we're told that the number of items in the two lists differs by 7. Any idea how we might write this as an equation?

OpenStudy (anonymous):

a+b-7=33?

OpenStudy (anonymous):

You're on the right track. Forget about the 33 for now. All we care about is that the first and second list differ by 7. So maybe the first list has 1 item and the second one has 8, or maybe the first list has 10 items and the second list has 17, or maybe the first list has 100 items and the second list has 107. If that makes sense, maybe try again with writing down an equation?

OpenStudy (anonymous):

(If that doesn't make sense, that's not a problem!)

OpenStudy (anonymous):

im not sure......A-7+B? what you say makes sense but im not sure about the equation

OpenStudy (anonymous):

Oooh, that's really close :) I think you might have mistyped something, cos an equation needs an equals sign!

OpenStudy (anonymous):

no i thought we were forgetting about 33 A-7+B= 33 ?

OpenStudy (anonymous):

Yep, we're ignoring the 33 for now. All we're trying to do is to turn this sentence into an equation: "The difference between a and b is 7."

OpenStudy (anonymous):

-7+a+b=33 i dont know the equation at all :(

OpenStudy (anonymous):

Okay, see if this makes sense: a - b = 7

OpenStudy (anonymous):

but i thought we were adding a + b?

OpenStudy (anonymous):

That was the first sentence. "The total number of items is 33" means: a + b = 33. Now we're looking at the second sentence: "The difference between the two numbers is 7" means: a - b = 7.

OpenStudy (anonymous):

okay i understand what now?

OpenStudy (anonymous):

Well, now we've got two equations (maybe you call them "simultaneous equations" at school?), and we'd like to solve them. Have you seen this kind of problem before?

OpenStudy (anonymous):

If not, I can walk you through it :)

OpenStudy (anonymous):

i dont think so? could you walk me through please??

OpenStudy (anonymous):

Okay, there are two ways to solve equations like these. One method is called "substitution" and the other is called "elimination". Have you seen either of these before? (If not, I'll decide which one to use!)

OpenStudy (anonymous):

i think substitution but im not sure

OpenStudy (anonymous):

Okay, there are two steps in this method. We'll start by taking the first equation ( a+b=33 ) and rearrange it so that one of the letters (let's say "a") is by itself on one side. Any idea how we might do this?

OpenStudy (anonymous):

no i dont think so..?

OpenStudy (anonymous):

Okay, right now we have "a" and "b" on the same side:. We need to get rid of the "b" from the left hand side.